I am trying to find the derivative of (1/x)
I need to see how to get passed the step of putting (x+h) in and dividing everything by h.
Any help is greatly apprieciated.

Answer by stanbon(75887)

I am trying to find the derivative of (1/x)
I need to see how to get passed the step of putting (x+h) in and dividing everything by h.
-----------------------
f(x+h) = 1/(x+h)
f(x) = 1/x
------
f(x+h)-f(x)=[1/(x+h) - 1/x] = [x-(x+h)]/[x(x+h)]
---
= [-h]/[x^2+hx]
-------------------------
Now divide by "h" to get:
= [-1]/[x^2+hx]
======================
Cheers,
Stan H.